OpenStudy (anonymous):

Brian has a clay vase shaped as a triangular prism with a base area of 45.5 square inches. It is one third filled with water. If there are 150 cubic inches of water in the vase, what is the height of the vase, rounded to the nearest tenth? 19.8 inches 9.9 inches 12.9 inches 6.3 inches

OpenStudy (anonymous):

I got 19.8

OpenStudy (anonymous):

Anyone?

OpenStudy (zzr0ck3r):

9.9 inches (1/3) V = 150 m^3

OpenStudy (anonymous):

Why is it 1/3 volume?

OpenStudy (zzr0ck3r):

"it is one third filled with water"

OpenStudy (anonymous):

Couldn't you multiply is by 3 to get it full?

OpenStudy (anonymous):

So it's 450?

OpenStudy (zzr0ck3r):

one sec I got to go move my car then ill explain

OpenStudy (anonymous):

Okay

OpenStudy (zzr0ck3r):

yes so V = 450 V = A*l 450 = A*l 450 = 45.5 L l = 450/45.5

OpenStudy (anonymous):

Don't you mean 45.5

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

then wouldn't you multiply by 2?

OpenStudy (anonymous):

Since it's a triangular prism?

OpenStudy (anonymous):

So the opposite of 1/2?

OpenStudy (anonymous):

is multiply by 2

OpenStudy (anonymous):

Hello?